# A Probabilistic Framework for Imputing Genetic Distances in Spatiotemporal Pathogen Models

### v2

@@ Abstract.
copyright: noneconference: SIGSPATIAL ’25, November 03–06, 2025, Minneapolis, MN, USA; ; $c$s: Applied computing Spatial epidemiologyccs: Computing methodologies Probabilistic reasoningccs: Applied computing Bioinformaticsccs: Information systems Spatial-temporal systems
@@ 4.2. Feature Construction
Before model training, all continuous variables are standardised to zero mean and unit variance. Categorical variables are expanded via one-hot encoding to enable compatibility with tree-based estimators. These transformations standardize the feature space and enable robust learning of complex interactions among epidemiologically relevant covariates.
@@ 7. Discussion and Conclusion
At higher divergence levels, predictive performance declines due to weaker correlations between metadata features and the true evolutionary distance. These limitations are expected in settings characterized by ecological discontinuities, long-range spread, or lineage reassortment, where genetic divergence may reflect independent introductions rather than continuous diffusion. Nonetheless, the model produces well-calibrated uncertainty intervals across a broad range of surveillance-relevant cases, demonstrating robustness in realistic outbreak scenarios.

### v3

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: The 33rd ACM International Conference on Advances in Geographic Information Systems; November 3–6, 2025; Minneapolis, MN, USAbooktitle: The 33rd ACM International Conference on Advances in Geographic Information Systems (SIGSPATIAL ’25), November 3–6, 2025, Minneapolis, MN, USAdoi: 10.1145/3748636.3762779isbn: 979-8-4007-2086-4/2025/11ccs: Applied computing Spatial epidemiologyccs: Computing methodologies Probabilistic reasoningccs: Applied computing Bioinformaticsccs: Information systems Spatial-temporal systems
@@ 4.2. Feature Construction
Before model training, all continuous variables are standardised to zero mean and unit variance. Categorical variables are expanded via one-hot encoding to enable compatibility with tree-based estimators. These transformations standardize the feature space and enable robust learning of complex interactions among epidemiologically relevant covariates.In addition to separate temporal and geographic predictors, we explicitly model their joint effect via the interaction term $\Delta t_{ij}\cdot d_{\mathrm{geo},ij}$ in order to capture spatiotemporal coupling, where rapid geographic spread over short intervals or prolonged temporal gaps at short distances may alter expected divergence.
For interpretability, we computed SHAP values for the median quantile model ($\tau{=}0.5$) using a tree-based explainer on the held-out test folds. We averaged absolute SHAP values per feature within each divergence stratum (LOW/MID/HIGH) and normalized them so importances sum to $100\%$ per stratum. The resulting shares are summarized in Fig. 3; per-feature values and scripts are provided in the repository.
@@ 6.4. Ablation Study
6.5. Feature Importance via SHAP
We computed SHAP values for the median ($\tau{=}0.5$) models and report normalized importances (share of total mean $|\mathrm{SHAP}|$) across divergence strata in Fig. 3. In the LOW stratum, geographic distance is the dominant driver, with temporal lag contributing secondarily, indicating strong spatial–temporal coupling when evolutionary distances are small. In MID, the relative influence of year increases alongside geographic distance and host-related indicators, consistent with gradual temporal drift and changing host composition within circulating lineages. In HIGH, year and host-related indicators account for a larger share of importance, while spatial/temporal predictors contribute less—aligning with scenarios where larger genetic gaps arise from independent introductions or lineage structure rather than continuous diffusion. These patterns mirror the ablation study (Table 3) and help explain why point accuracy declines in HIGH while interval calibration remains stable.
Figure 3. Normalized SHAP importance for the median model ($\tau{=}0.5$) across divergence strata.
@@ 7. Discussion and Conclusion
Although demonstrated on highly pathogenic avian influenza A/H5, the framework is pathogen-agnostic. Any setting with rich pairwise metadata including temporal lag, spatial separation, host or demographic attributes and observed divergence for a subset of pairs can be used. Adapting to other pathogens would require selecting an appropriate evolutionary substitution model and adjusting temporal scales to match evolutionary rates.
At higher divergence levels, predictive performance declines due to weaker correlations between the available metadata features and the true evolutionary distance. This effect is most pronounced in cases where divergence is driven by between-lineage differences, reassortment events, or other major genomic changes that are not explicitly represented in the model’s feature set. Without lineage or subclade identifiers, the model relies solely on spatiotemporal and host covariates, which have limited explanatory power in such contexts. These limitations are expected in settings characterised by ecological discontinuities, long-range spread, or independent introductions into distinct host populations, where genetic divergence may not follow patterns of continuous spatial diffusion. Consistent with this interpretation, the SHAP summaries for the median model (Fig. 3) show a shift in importance away from spatial and temporal predictors and toward year and host-related indicators in the HIGH stratum, mirroring the ablation results (Table 3). While this constrains accuracy in the HIGH divergence stratum, the model nonetheless produces well-calibrated uncertainty intervals across a broad range of surveillance-relevant cases, demonstrating robustness in realistic outbreak scenarios.